RhondaSommer (rhondasommer):

Given 🔺ABC with a = 24 b=13.75, angle A = 108⁰ and angle B = 33⁰, find c. Round the sine value to the nearest thousandth and the answer to the nearest hundredth. 39.00 17.83 15.87 13.74

RhondaSommer (rhondasommer):

Sean :)

OpenStudy (sweetburger):

Angle C is found easily by 180 - 108 -33 = 39

RhondaSommer (rhondasommer):

 hmm ok

OpenStudy (sweetburger):

So we use the law of sins again because it states \[\sin(c)/c = \sin(a)/a = \sin (b) /b\] so we will only consider \[\sin(a)/a = \sin(c)/c\]

OpenStudy (sweetburger):

\[\sin(108)/(24) = \sin(39)/c \] solve for c

RhondaSommer (rhondasommer):

0.951*C=24*0.629 0.951*C=15.096 C=15.87

OpenStudy (sweetburger):

looks good @RhondaSommer

RhondaSommer (rhondasommer):

sweer

RhondaSommer (rhondasommer):

Given 🔺ABC with a = 9 angle A = 80⁰ and angle B = 35⁰, find b. Round the sine value to the nearest thousandth and the answer to the nearest hundredth. 5.24 6.53 8.28 67.00